Hints and Tips for Renovating an early Tektronix 475: Timebase, Switches, Potentiometers, and HF Response

I recently picked up an early Tek 475 scope plus service manual plus probes at a hamfest. I asked what was wrong, and the seller indicated the timebase was unreliable. Fair enough and, given the low price, at worst it would be a parts mule.

It was a bit mucky on the outside, as would be expected of something that had been in storage. A quick look the scope showed it was cleaner inside than I expected, and playing with it made me feel it was worth resuscitating. If nothing else, the puzzles involved would be more interesting and worthwhile than crosswords and soduku.

Overall the renovation involved some traditional problems and solutions, plus a couple that are less well-known – and a few diversions down blind alleys. Maybe the hints and tricks outlined below will save other people some time.

The problems were:

  • fan not working
  • timebase working but unreliable
  • both vertical channels working but unreliable
  • unreliable noisy pots
  • mechanically sticky and unreliable switches


All ranges and the A/B sweeps worked, but not every time. Jiggling the timebase knob vigorously helped, eventually. The solution to this was a bit of a pig, and I haven’t seen it referred to elsewhere.

The principal problem was related to the microswitch which detects whether the B timebase is locked to the A timebase, or is on a faster sweep speed. The timebase knob contains a pin which rests on an asymmetric coaxial white plastic lever, which rests on the black microswitch (on the right in the pictures). The knob pushes the pin in one direction, and thereby pushes the microswitch lever down. The force in the other direction is provided by the microswitch, not by a dedicated spring – an irritatingly cheap and nasty design. That force was insufficient to reliably move the pin.

The root cause was that the pin was sticking slightly inside the knob, for reasons unknown. There was no room to add a spring, so I considered using a drop of light machine oil, but decided it would probably be too viscous for the weak force from the microswitch.

Disassembling the mechanism and repeatedly rinsing with IPA solved that problem. Good, but it did raise the spectre of invisible gunk on contacts elsewhere in the scope.

I cleaned the timebase’s cam-operated gold finger contacts using the traditional piece of smooth paper soaked in IPA. That was a significant improvement, but not a complete cure. In the end I added a tiny drop of Caig DeOxit to each finger, using the smallest jewellers screwdriver available. I’d have preferred not to do that since it might trap dust over time. It also reinforces the possibility of gunk.

The timebase ranges are now usable, but sometimes the knob has to be lightly jiggled into position.

Due to the limited space, reinserting the timebase assembly is an interference fit! It is easier if you remove the assembly, pop out the small white front-panel plastic bezels, re-insert the assembly, and finally pop the bezels back in.

Vertical Sensitivity Controls

These were also a pig, but if I had realised how bits were held together, it would have been significantly easier.

The principal problem was that some of the HF compensations were grossly inaccurate.

The 2mV and 5mV sensitivity settings weren’t bad. The sensitivity and offsets were a little out, but that is easily tweaked following the manual. The HF response was pretty good, but following the manual allowed very minor tweaks to produce a clean 1.8ns risetime. That’s good since it indicates most of the scope is working well.

The other sensitivity settings, related to the 2/4/10/100 attenuators were the problem. Tweaking the compensation capacitors for the 10mV, 20mV, 50mV, 500mV ranges (/2, /4, /10, /100 attenuators) was easy, as expected. But having two attenuator sections in series for the 100mV, 200mV, 1V, 2V, 5V ranges showed significant errors. That’s a problem since they are supposed to be correct and there is nothing to be tweaked.

Buzzing the attenuator’s cam-operated gold finger contacts indicated decent DC continuity. Swapping in some attenuators from a borrowed 475 significantly improved matters, but was ambiguous since I did not want to alter the compensation capacitor’s settings. That was sufficient to indicate the scope was generally good, but not to determine whether the attenuators were faulty (a traditional problem) or there was gunk on the contacts and traces, similar to the A/B timebase pin.

Some of the capacitors were extraordinarily sensitive to their position, in a way suggesting a poor contact. I decided to carefully re-solder the mounting pin connections inside each attenuator, since some have reported that can improve things. It did improve the sensitivity, but did not resolve the principal HF compensation problem.

I decided to clean the contacts. Since some are buried in the switch, access requires removal of the preamp board followed by the attenuator and switch assembly – or so I thought.

I removed the vertical preamp board, but did not dare remove the coax lines for fear of damaging the stiff coax or petolta connection. I left the board flapping but loosely supported. Undesirable, but sufficient.

The attenuator and channel switch assembly were then removed. The four 1/4 inch nuts that secure it to the front panel are a pig. There’s no way a spanner can reach them, I don’t have a 1/4 inch nut runner/spinner, and neither do any of the local shops. Fortunately I realised that extension shanks for bog-standard screwdriver bit sets are 1/4 inch; they just fitted in the space available and the nuts were removed.

1/4″ Nuts Behind Capacitor, Hex Standoff Foreground Left

After removing the attenuator and switch assembly it became clear that it would have been sufficient to remove only the 1/4 inch nuts plus the hex standoffs between the attenuator PCB and attenuator shield. The preamp board and switch assembly could have been left untouched. Oh well, live and learn.

Unfortunately cleaning contacts did not improve things. Time to ask the TekScopes group.

The principal contribution, from Fabio Trevisan on TekScopes, was to note that adjusting the attenuators requires a “normaliser”. I had been following the service manual and using

  • a 2.5V 50Ω step generator with <1ns risetime
  • an RF attenuator to reduce the amplitude
  • a 50Ω through terminator attached directly to the scope input

That gives the correct amplitude and risetime, but the 50Ω gives invalid time constants when driving the 1MΩ attenuator sections. A normaliser mimics the time constants of a *10 probe, and consists of a 1MΩ//20pF in the signal line between two BNC sockets. The normaliser is inserted between the 50Ω through terminator and the scope’s input.

I quickly threw the relevant components together and there was a significant improvement – but it was still far from perfect. Eventually I noticed that the corner edge was better/worse depending on the pulse’s amplitude, and that was the case within a single range and on a known good scope. Now I had used a 20pF C0G capacitor, which is supposedly not voltage and temperature sensitive; either I have a counterfeit component or there is a noticeable voltage sensitivity. Replacing it with an air dielectric trimcap removed the problem.

Finally I was in a position to tweak the attenuator’s capacitors. After a few loops this was achieved, resulting in a nice 1.8ns risetime on all vertical sensitivities and with *10 probes or a direct  50Ω terminations.

Having said that, some of the attenuator trimcaps are still give the impression there might be a poor contact. Nonetheless after dropping the scope an inch in various axes, it still works. Time to call that a day, and move on.

Sticky Switches

Some of the front panel switches made unreliable electric contact, and some were reluctant to mechanically return to their normal position without assistance.

The standard multipole PCB switches (e.g. vertical and horizontal selection) were lubricated and cleaned using Caig DeOxit. That resolved the electrical problems.

Mechanically the non-latching switches were fine, but the latching switches were reluctant to pop out. The problem was that the small pin under the bronze spring was sticking rather than jiggling around. A small drop of light machine oil solved that, and all switches work well.

The bandwidth control’s cam-operated gold finger PCB contacts were cleaned using paper and IPA, wherever there was sufficient access. That control’s brown plastic housing was cleaned similarly as far as possible, but the bandwidth switch remained mechanically sticky. I added a small drop of light machine oil to the sliding brown plastic “cam” housing. I’d have preferred not to do that since over time the oil will retain dust and dry out, but the controls are now usable.

U/S Fan

This was easy. I simply replaced the quad transistor array motor driver with four ZTX630 transistors that I had in my junk box. They have a better h_{FE}, I_{cmax} and P_{d} than the original, but obviously they are neither matched nor thermally coupled.

I don’t know why the original has the array. It seems to work well enough as it is, so I haven’t bothered to attempt to thermally couple the ZTX630s.

Unreliable Potentiometers

A traditional problem. The potentiometers on the front panel are easy to replace, those mounted directly on the PCBs less so.

All I did was to loosen as many of the pot’s securing bolts as possible, then drip in some Caig DeOxit FaderLube and re-tighten the bolts. I avoided pots with HV across them, for obvious reasons.

The pots are significantly improved. They are now usable and operate smoothly when rotated slowly, but are a bit noisy when rotated quickly. How usable they will be in the long-term remains to be seen.

Leave a Reply

Fill in your details below or click an icon to log in:

WordPress.com Logo

You are commenting using your WordPress.com account. Log Out /  Change )

Google photo

You are commenting using your Google account. Log Out /  Change )

Twitter picture

You are commenting using your Twitter account. Log Out /  Change )

Facebook photo

You are commenting using your Facebook account. Log Out /  Change )

Connecting to %s